An object has a mass of 200 g and a volume of 125 mL.  What is its density in g/mL?


1
Expert's answer
2021-03-24T19:44:02-0400
"\rho=\dfrac{m}{V}=\dfrac{200\ g}{125\ mL}=1.6\ \dfrac{g}{mL}."